OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 108, -18, 3, -1/2

OpenStudy (anonymous):

it explodes

jimthompson5910 (jim_thompson5910):

what type of sequence is this? any ideas?

OpenStudy (anonymous):

arithmetic?

jimthompson5910 (jim_thompson5910):

no it's not arithmetic

OpenStudy (anonymous):

then geometric.

OpenStudy (anonymous):

you sere its not power

jimthompson5910 (jim_thompson5910):

subtract any two consecutive terms and you'll find the differences aren't the same so there's no common difference

jimthompson5910 (jim_thompson5910):

yes it's a geometric sequence

jimthompson5910 (jim_thompson5910):

first term is a = 108 common ratio is r= -18/108 = -1/6

OpenStudy (anonymous):

I see, is that a way to always prove its not arithmetic?

OpenStudy (anonymous):

so we're going to use a formula for this right?

OpenStudy (anonymous):

this jim is helping so good

OpenStudy (anonymous):

hey jim you know anything about transport phenomena

OpenStudy (anonymous):

What do i do with the "a" and "r". Is there a formula?

OpenStudy (anonymous):

jimbo you know anything about dimensional analysis.

OpenStudy (anonymous):

Type geometric series into google and it will give you an easy reference

OpenStudy (anonymous):

Try Pauls Online Math notes but this kinda stuff should be in your book.

jimthompson5910 (jim_thompson5910):

sry the site is going slow for me but you then use the formula S = a/(1 - r) to find out what it converges to